OpenStudy (anonymous):

Solve 2x2 - 8x = -7 using the quadratic formula?

OpenStudy (anonymous):

you have to use a better notation. is it 2x^2-8x=-7 ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okey so 2x^2-8x+7=0 right? The general formula for this kind of equations is ax^2+bx+c so a=2 b=-8 and c=7 use the quadratic formula .

OpenStudy (anonymous):

Could you work that out and tell me what you did, using the formula? I just started learning this today so I am still a bit confused!

OpenStudy (anonymous):

yes, look at the pic. ask the parts you dont understand. after finding a b and c you just plug in the values. hope you can see it clearly my handwriting is not decent :D

OpenStudy (anonymous):

Hold on, taking a long time to load the picture :P

OpenStudy (anonymous):

Can I have a download link? It doesn't want to open >.<

OpenStudy (anonymous):

try this

OpenStudy (anonymous):

Ok thats better!

OpenStudy (anonymous):

my answer is the same dude :D just different notations.